how do you graph the equation:
y-3=-3(x+1)

Answer by Alan3354(69443)

It's a straight line, so you need 2 points.
y-3=-3(x+1)
y-3=-3x - 3
y = -3x
Pick 2 values for x and find y
Plot the 2 points, draw a line thru them.